# CorMulT: A Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

### v2

@@ Abstract
Multimodal sentiment analysis is an active research area that combines multiple data modalities, e.g., text, image and audio, to analyze human emotions, and benefits a variety of applications. Existing multimodal sentiment analysis methods can be roughly classified as modality interaction-based methods, modality transformation-based methods and modality similarity-based methods. However, most of these methods highly rely on the strong correlations between modalities, and cannot fully uncover and utilize the correlations between modalities to enhance sentiment analysis. Therefore, these methods usually achieve unsatisfactory performance for identifying the sentiment of multimodal data with weak correlations. To address this issue, we proposed a two-stage semi-supervised model termed Correlation-aware Multimodal Transformer (CorMulT) which consists of pre-training stage and prediction stage. At the pre-training stage, a modality correlation contrastive learning module is designed to efficiently learn modality correlation coefficients between different modalities. At the prediction stage, the learned correlation coefficients are fused with modality representations to make the sentiment prediction. According to the experiments on the popular multimodal dataset CMU-MOSEI, CorMulT obviously surpasses the state-of-the-art multimodal sentiment analysis methods.
@@ IV-A1 Dataset
We conduct experiments over the CMU_MOSEI dataset which is a widely used benchmark for multimodal sentiment analysis. This dataset contains 3,228 distinct video clips that cover diverse domains, including reviews, debates and consults. Each video clip has three modalities, i.e., video, audio and text, and is labelled with seven classes of fundamental emotions with an intensity scale from -3 to 3. Within these clips, there are a total of 23,453 sentences, with each sentence along with its corresponding audio and video segments jointly constituting a sample. Table I provides the statistical details of CMU_MOSEI dataset. For our experiments, we allocate 70% of the dataset(16,417 sentences) to the training set, 15% to the validation set(3,518 sentences), and 15% to the test set(3,518 sentences).
@@ IV-A2 Experiment Settings
All experiments were conducted on a server equipped with an Intel SkyLake-E 12-core processor and a GeForce GTX 1180 graphics card. We run each experiment five times, and report the average results. Multiple evaluation metrics as listed below are introduced to evaluate the performance of CorMult model and the baseline multimodal sentiment analysis methods.
@@ IV-B Effects of Perturbation Strategy
Each strategy is applied across the entire training dataset. Training parameters and model architectures are consistent across strategies to ensure a fair comparison. Table II presents the sentiment analysis results while using different data perturbation strategies. Among the four strategies, Strategy D achieves the highest values in terms of all five performance evaluation metrics, indicating the advantage of combining multiple data perturbation strategies.
TABLE II: The results of CorMulT when using different data perturbation strategies for contrastive learning in MCE module.
@@ IV-C Effects of Distance Measurement
Table III presents the results of CorMulT while using different distance measurement methods. According to the results, cosine similarity emerges as the optimal distance metric for MCE, and could efficiently capture the inter-modal correlations.
TABLE III: Results of CorMulT while using different distance measurement methods.
@@ IV-E Model Comparison
CorMulT model is compared with multiple representative multimodal sentiment analysis models, including EF-LSTM[32], LMF[33], TFN[34], MFM[35], CM-Bert[36], MISA[8], BBFN[4], MULT[18], ICCN[37] and MMLatch[5]. Table IV presents the results of CorMulT and baseline methods, where V+A+T indicates that the model utilizes vision, audio and text modalities. CM-Bert only uses vision and text modalities, i.e., V+T. Among all models, CorMulT exhibits superior performance. Specifically, considering the F1 score and accuracy metrics (Acc@2 and Acc@7), CorMulT distinctly outperforms other models. For Acc@7, CorMulT achieves 52.8%, surpassing the next best, i.e., MMLatch, at 52.1%. Similarly, for Acc@2, CorMulT scores 86.5%, notably better than MISA’s 83.9%. Regarding the F1 score, CorMulT’s 86.7% markedly eclipsed others, showcasing its robustness since F1 is the harmonic mean of precision and recall. Lastly, examining Corr (correlation), CorMulT excels with a score of 0.779, substantially ahead of BBFN’s 0.761, underscoring its adeptness in comprehending multimodal data intricacies. In sum, CorMulT model manifests significant superiority in multimodal sentiment analysis in terms of accuracy, F1 scores, and correlation.
TABLE IV: Comparison of CorMulT with multiple baseline models for multimodal sentiment analysis.
@@
EF-LSTM V+A+T 43.8 75.2 75.3 0.609
LMF V+A+T 48.1 82.1 82 0.677
TFN V+A+T 50.2 82.5 82.1 0.701
MFM V+A+T 50.3 82.4 82.3 0.717
CM-Bert V+T 50.5 83.6 83.5 0.747
MISA V+A+T 50.7 83.9 83.8 0.748
BBFN V+A+T 50.8 82.2 82.2 0.761
MULT V+A+T 51.1 83.3 83.5 0.758
ICCN V+A+T 51.2 82.2 82.4 0.713
MMLatch V+A+T 52.1 83.3 83.9 0.754
CorMulT V+A+T 52.8 86.5 86.7 0.779
IV-F Ablation Studies
@@
For modality-specific feature extraction, we applied tailored approaches to address the distinct characteristics of three different data types. To gauge the effectiveness of these approaches, we specifically examined the transformation of the Mel spectrum matrix for audio data and the application of ResNet50 for image feature mapping. Table VI quantitatively evaluates the effectiveness of different feature extraction strategies in the MCE module across various modalities. The strategies include the use of Mel spectrum transformation for audio data, ResNet50 for image data, both methods combined, and a baseline with no specialized feature extraction methods. Metrics assessed include loss values for audio ($Loss_{A}$), text ($Loss_{T}$), and visual ($Loss_{V}$) data, along with the average loss ($Loss_{ave}$) and epoch time ($t_{epoch}$). The results demonstrate the efficacy of modality-specific feature extraction. Tailored approaches such as Mel spectrum for audio and ResNet50 for visual data significantly reduce the losses across respective modalities, with the combined approach yielding the lowest average loss ($Loss_{ave}$ = 0.168) and shortest training time (41.02 seconds per epoch). This verifies the importance of employing specific feature extraction methods for each modality.
TABLE V: The results of CorMulT while using different modality-specific feature extraction methods.
@@
We also assessed CorMulT’s performance with different combinations of pre-trained modalities, including no modality specified (None), audio and text only (A+T), visual and text only (V+T), and visual, audio, and text (V+A+T). The results shown in Table VI demonstrate that CorMulT’s performance is enhanced by incorporating more modalities during pre-training. The results underscore the significance of inter-modality synergy in improving sentiment analysis outcomes. This matches the expectation that the more modality similarity information that CorMulT learns, the higher performance it achieves.
TABLE VI: The results of CorMulT while using different modalities for pre-training.

### v3

# CorMulT: A Semi-supervised Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

@@ Abstract
Multimodal sentiment analysis is an active research area that combines multiple data modalities, e.g., text, image and audio, to analyze human emotions, and benefits a variety of applications. Existing multimodal sentiment analysis methods can be roughly classified as modality interaction-based methods, modality transformation-based methods and modality similarity-based methods. However, most of these methods highly rely on the strong correlations between modalities, and cannot fully uncover and utilize the correlations between modalities to enhance sentiment analysis. Therefore, these methods usually achieve unsatisfactory performance for identifying the sentiment of multimodal data with weak correlations. To address this issue, we proposed a two-stage semi-supervised model termed Correlation-aware Multimodal Transformer (CorMulT) which consists of pre-training stage and prediction stage. At the pre-training stage, a modality correlation contrastive learning module is designed to efficiently learn modality correlation coefficients between different modalities. At the prediction stage, the learned correlation coefficients are fused with modality representations to make the sentiment prediction. According to the experiments on the popular multimodal dataset CMU-MOSEI, CorMulT obviously surpasses the state-of-the-art multimodal sentiment analysis methods. The code of CorMulT is available at https://github.com/YAMY1234/CorMulT.
@@ IV-A1 Dataset
We conduct experiments primarily on the CMU_MOSEI dataset, a widely used benchmark for multimodal sentiment analysis, and supplement our study with additional validation on the CH-SIMS dataset to assess model generalizability across different datasets.
The CMU_MOSEI dataset comprises 3,228 distinct video clips spanning various domains such as reviews, debates, and consultations. Each clip contains three modalities—video, audio, and text—and is annotated with seven fundamental emotion classes on an intensity scale ranging from -3 to 3. In total, the dataset includes 23,453 sentences, where each sentence, along with its corresponding audio and video segments, forms a sample. Table II presents the statistical details of the CMU_MOSEI dataset. For our experiments, we split the dataset into 70
To further evaluate the model’s performance on a different dataset, we also conduct experiments on CH-SIMS, a newly established benchmark for multimodal sentiment analysis in Chinese. The CH-SIMS dataset consists of 2,281 curated video segments sourced from movies, TV series, and variety shows. Each segment is annotated with both multimodal and unimodal sentiment labels, facilitating a detailed analysis of modality-specific contributions. The dataset covers three modalities—text, visual, and audio—and sentiment labels are categorized into five levels: Negative, Weakly Negative, Neutral, Weakly Positive, and Positive. Table II provides the detailed statistics of the CH-SIMS dataset.
This setup allows us to conduct a comprehensive set of comparisons and ablation studies on CMU_MOSEI while verifying the model’s robustness and effectiveness on CH-SIMS.
@@
TABLE II: The overview of CH-SIMS dataset.
Feature Description
Source Movies, TV series, Variety shows
Total Samples 2,281
Video Segments 2,281
Frame Rate 30Hz
Unique Speakers 474
Average Segment Length 3.67 seconds
Average Word Count 15 words
Modalities Text (T), Visual (V), Audio (A)
Sentiment Annotations Multimodal and Independent Unimodal
Labels Three categories: Negative,Neutral, Positive
@@ IV-A2 Experiment Settings
All experiments were conducted on a server equipped with two Intel SkyLake-E 12-core processor and a GeForce GTX 2080 graphics card. We run each experiment five times, and report the average results. Multiple evaluation metrics as listed below are introduced to evaluate the performance of CorMult model and the baseline multimodal sentiment analysis methods.
@@ IV-B Effects of Perturbation Strategy
Each strategy is applied across the entire training dataset. Training parameters and model architectures are consistent across strategies to ensure a fair comparison. Table III presents the sentiment analysis results while using different data perturbation strategies. Among the four strategies, Strategy D achieves the highest values in terms of all five performance evaluation metrics, indicating the advantage of combining multiple data perturbation strategies.
TABLE III: The results of CorMulT when using different data perturbation strategies for contrastive learning in MCE module.
@@ IV-C Effects of Distance Measurement
Table IV presents the results of CorMulT while using different distance measurement methods. According to the results, cosine similarity emerges as the optimal distance metric for MCE, and could efficiently capture the inter-modal correlations.
TABLE IV: Results of CorMulT while using different distance measurement methods.
@@ IV-E Model Comparison
On CMU_MOSEI dataset, CorMulT model is compared with multiple representative multimodal sentiment analysis models, including EF-LSTM[32], LMF[33], TFN[34], MFM[35], CM-Bert[36], MISA[8], BBFN[4], MULT[18], ICCN[37], MMLatch[5], TMSON[38], DVMI[39], ASBA-TLRF[40], EMT[41] and MMIN[42]. Table V presents the results of CorMulT and baseline methods, where V+A+T indicates that the model utilizes vision, audio and text modalities. CM-Bert only uses vision and text modalities, i.e., V+T. Among all models, CorMulT exhibits superior performance. Specifically, considering the F1 score and accuracy metrics (Acc@2 and Acc@7), CorMulT distinctly outperforms other models. For Acc@7, CorMulT achieves 52.8%, surpassing the next best, i.e., MMLatch, at 52.1%. Similarly, for Acc@2, CorMulT scores 86.8%, surpassing the next best, i.e., DVMI, at 86.6%. Regarding the F1 score, CorMulT’s 86.7% markedly eclipsed others, showcasing its robustness since F1 is the harmonic mean of precision and recall. Lastly, examining Corr (correlation), CorMulT excels with a score of 0.779, substantially ahead of EMT’s 0.774, underscoring its adeptness in comprehending multimodal data intricacies.
On CH-SMIS dataset, CorMulT model is compared with multiple representative multimodal sentiment analysis models, including EF-LSTM[32], LF-DNN[43], TFN[34], LMF[33], MFM[35], Graph-MFN[44], MULT[18], MLF-DNN[45], MTFN[45], Self-MM[46], DVMI[39], M3SA[47] and EMT[41]. Table VI presents the results of CorMulT and baseline methods. Specifically, considering the F1 score and accuracy metrics (Acc@2 and Acc@3), CorMulT distinctly outperforms most of other models. For Acc@2, CorMulT achieves 83.4%, surpassing the next best, i.e., MTFN, at 82.5%. Similarly, for the F1 score, CorMulT’s 83.5% markedly eclipsed others, showcasing its robustness for binary sentiment classification. Besides, CorMulT excels with a score of 0.691 in Corr(Correlation), substantially ahead of MLF-DNN’s 0.675, underscoring its adeptness in comprehending multimodal data intricacies. Regarding Acc@3, CorMulT scores 68.0%, maintaining its strong performance but lower than MLF-DNN’s 69.0. By examining the results, we observe that our model performs well in distinguishing between positive and negative sentiments. However, seemingly due to the perturbations introduced during training, it tends to classify more ambiguous emotions as neutral (as shown in our results that there are more non-neutral samples classified as neutral). This contributes to a slightly lower performance in the triple class classification task. Nevertheless, CorMulT remains among the top-performing models in this aspect compared to others.
In sum, CorMulT model manifests significant superiority in multimodal sentiment analysis in terms of accuracy, F1 scores, and correlation.
TABLE V: Comparison of CorMulT with multiple baseline models for multimodal sentiment analysis on CMU_ MOSEI dataset.
@@
EF-LSTM(2018)[32] V+A+T 43.8 75.2 75.3 0.609
LMF(2018)[33] V+A+T 48.1 82.1 82 0.677
TFN(2017)[34] V+A+T 50.2 82.5 82.1 0.701
MFM(2023)[35] V+A+T 50.3 82.4 82.3 0.717
CM-Bert(2020)[36] V+T 50.5 83.6 83.5 0.747
MISA(2020)[8] V+A+T 50.7 83.9 83.8 0.748
BBFN(2021)[4] V+A+T 50.8 82.2 82.2 0.761
MULT(2019)[18] V+A+T 51.1 83.3 83.5 0.758
ICCN(2019)[37] V+A+T 51.2 82.2 82.4 0.713
MMLatch(2022)[5] V+A+T 52.1 83.3 83.9 0.754
TMSON(2024)[38] V+A+T 55.6 86.4 86.2 0.766
DVMI(2024)[39] V+A+T - 86.6 86.4 0.773
ABSA-TLRF(2024)[40] V+A+T 54.8 86.5 86.6 -
EMT(2023)[41] V+A+T 54.5 86.0 86.0 0.774
MMIN(2024)[42] V+A+T - 85.9 85.8 0.761
CorMulT V+A+T 55.9 86.8 86.7 0.779
TABLE VI: Comparison of CorMulT with multiple baseline models for multimodal sentiment analysis on CH-SIMS dataset.
Model Modality Acc@2 Acc@3 F1 Corr
EF-LSTM(2018)[32] V+A+T 69.4 54.3 81.9 0.055
LF-DNN(2013)[43] V+A+T 79.0 64.3 80.2 0.555
TFN(2017)[34] V+A+T 80.7 65.1 81.6 0.591
LMF(2018)[33] V+A+T 79.2 64.7 81.6 0.576
MFM(2018)[48] V+A+T 77.9 65.7 79.9 0.582
Graph-MFN(2018)[44] V+A+T 78.8 65.7 78.2 0.578
MulT(2019)[18] V+A+T 78.6 64.8 79.7 0.564
MLF-DNN(2020)[45] V+A+T 82.3 69.0 82.5 0.675
MTFN(2020)[45] V+A+T 82.5 68.8 82.6 0.670
Self-MM(2021)[46] V+A+T 80.0 65.5 80.4 0.592
DVMI(2024)[39] V+A+T 81.2 - 81.3 0.599
M3SA(2024)[47] V+A+T 81.6 68.1 81.5 0.614
EMT(2023)[41] V+A+T 80.1 67.4 80.1 0.623
CorMulT V+A+T 83.4 68.0 83.5 0.691
IV-F Visualization Analysis
To better demonstrate how the proposed Modality Correlation Evaluation (MCE) module captures and distinguishes modality relationships, we visualize in Fig. 5 the final-layer representations (after MCE) for three positive samples and three negative samples. Each subfigure displays a heatmap for one modality (Text, Audio, or Vision), where the vertical axis corresponds to time and the horizontal axis represents the feature dimension. The color intensity indicates the learned feature values; redder tones reflect higher feature values, whereas bluer tones indicate lower feature values. In this example, we employ our perturbation “strategy 3” by injecting Gaussian noise into the Vision modality to create negative samples.
Positive vs. Negative Samples. In the top row of Fig. 5, we show the learned representations for the three modalities when they are highly correlated (positive samples), while the bottom row demonstrates the outcome for negatively perturbed samples (negative samples). As shown in the figure, the Vision representations in the negative case appear more misaligned and deviate substantially from the other two modalities (Text and Audio).
On the right side of Fig. 5, we also include quantitative similarity scores measured by cosine similarity. One can observe that for positive samples, the cross-modal similarities (Text–Audio, Text–Vision, and Audio–Vision) remain consistently high (around 0.60.60.60.60.70.70.70.7). By contrast, in the negative samples, only the Text–Audio similarity stays at a relatively higher value (around 0.60.60.60.60.70.70.70.7), while both Text–Vision and Audio–Vision drop to $\approx 0.4$ or lower. This indicates that the model correctly identifies the perturbation in the Vision modality and registers much lower correlation between Vision and the other two modalities.
Figure 5: Visualization of the final-layer MCE features for three positive samples (top row) and three negatively perturbed samples (bottom row). The color bar on the right indicates the feature values, with redder tones signifying higher activation. On the far right, the corresponding pairwise cross-modal similarities are shown, illustrating strong positive correlations in the top row and reduced Vision correlations in the negative samples.
Difference Maps. To further highlight how perturbations affect modality relationships, we plot in Fig. 6 the pairwise difference heatmaps between modalities. For instance, Text–Audio (pos) displays the difference between the Text and Audio features in a positive sample, while Text–Vision (neg) shows the difference for a negative sample where Vision has been perturbed. As the color scale indicates, deeper red/blue patches imply larger absolute differences. We observe that for positive samples, the differences between any two modalities remain relatively small, reflecting their higher correlation. In contrast, the negative examples exhibit markedly higher difference values in Text–Vision and Audio–Vision, reinforcing the fact that these modalities have become misaligned due to the Gaussian noise applied to Vision.
Figure 6: Heatmaps of pairwise feature differences between modalities, with the top row corresponding to positive samples and the bottom row to negative samples. Larger red/blue patches correspond to bigger differences. Notice that in the negative samples, the difference maps (especially for Vision) exhibit much larger intensities.
These visualizations confirm that our proposed MCE module indeed captures modality correlations effectively. When modal inputs are properly aligned and correlated (positive samples), the learned representations show relatively small differences and high cross-modal similarity. However, once a modality (Vision in this case) is perturbed or not correlated as negative samples, the MCE outputs reflect comparably lower similarity with the unperturbed modalities. This demonstrates the module’s capability to discern correlated vs. uncorrelated modalities, thereby enhancing the overall transformer architecture’s ability to learn robust multimodal representations.
IV-G Weak Modality Correlation Analysis
This section provides targeted experiments to evaluate model performance on four types of weak correlations and compares results with both strongly correlated and baseline cases, i.e., voice discrepancy, content discrepancy, alignment discrepancy, and clarity discrepancy.
These discrepancies are tied to three negative-sample perturbation strategies in our contrastive learning module: Strategy A randomly replaces one modality with another sample (e.g., generating voice/content discrepancy), Strategy B shifts the temporal alignment among modalities (alignment discrepancy), and Strategy C injects noise or random corruption (clarity discrepancy). Specifically, these scenarios are simulated with the negative sample generation strategies by:
  • Voice Discrepancy: Audio content is replaced or shifted, leading to mismatched semantics with the text/visual streams.
  • Content Discrepancy: Text or visual frames are replaced by other samples, inducing semantic inconsistencies.
  • Alignment Discrepancy: Audio, visual, or textual elements are temporally offset, introducing misalignment among modalities.
  • Clarity Discrepancy: Noise or blurring is added to audio/visual modalities, degrading signal quality.
Comparison with Strongly Correlated Data
Table VII first shows that on the strongly correlated portion of CMU_MOSEI, our CorMulT achieves $55.9\%$ on Acc@7 and $86.7\%$ on F1. Once the dataset is disturbed with random weak-correlation perturbations, the performance drops to $51.7\%$ (Acc@7) and $82.3\%$ (F1), but still substantially outperforms the baseline MULT [18] under the same perturbations (Acc@7: $47.5\%$, F1: $76.3\%$). This validates that our modal-correlation contrastive learning framework effectively mitigates the adverse impacts of weak correlation.
TABLE VII: CorMulT vs. MULT on strongly vs. weakly correlated data (overall random perturbations).
Method Data Type Acc@7 F1 Corr
CorMulT Strongly Corr. 55.9 86.7 0.779
CorMulT Weakly Corr. 51.7 82.3 0.735
MULT Weakly Corr. 47.5 76.3 0.635
Analysis of Four Weak-Correlation Types
To systematically evaluate our model’s robustness against different forms of weak modality correlation, we design four controlled perturbation scenarios, each reflected by a distinct negative-sample generation strategy: (Strategy A) random replacement for voice/content discrepancy, (Strategy B) temporal shifting for alignment discrepancy, and (Strategy C) noise corruption for clarity discrepancy. Table VIII reports the comparison between our CorMulT and a standard multimodal transformer MULT under each discrepancy type. We observe:
  • Content/Voice Discrepancy (Strategy A). By replacing the text or visual modality (content discrepancy) or substituting the modality (voice discrepancy), we induce comparable semantic conflicts across modalities. As expected, both models suffer a notable performance drop. Nonetheless, CorMulT still benefits from partially correlated cues in the unaffected modalities, retaining mid-70% F1 and consistently outperforming MULT in accuracy, F1, and correlation.
  • Alignment Discrepancy (Strategy B). Time misalignment is introduced by shifting the timeline of one modality. CorMulT’s correlation-aware design and contrastive objectives help mitigate the negative effects of misalignment more effectively, preserving higher accuracy and stronger correlation compared to MULT.
  • Clarity Discrepancy (Strategy C). With Strategy C, we degrade one modality’s correlation and quality via noise addition or blurring. Despite the corruption, CorMulT exhibits relatively stable performance, suggesting robustness to signal noise. This contrasts with the more pronounced degradation observed in MULT, indicating that CorMulT’s contrastive approach better preserves essential features under noise perturbation.
TABLE VIII: Comparison of CorMulT and MULT under different weak-correlation types. Each discrepancy is simulated by a distinct negative-sample generation strategy (A/B/C).
Discrepancy Method Acc@7 (%) Acc@2 (%) F1 (%) Corr
Content (A) CorMulT 44.2 74.4 74.4 0.639
MULT 41.5 70.2 70.5 0.598
Voice (A) CorMulT 47.6 76.8 76.8 0.657
MULT 45.1 73.8 73.6 0.623
Alignment (B) CorMulT 53.3 85.2 85.2 0.752
MULT 48.7 78.5 78.8 0.690
Clarity (C) CorMulT 53.6 84.0 84.9 0.749
MULT 48.2 77.8 78.1 0.682
Conclusion on Weak-Correlation Analysis. Our targeted experiments demonstrate that CorMulT effectively mitigates the adverse effects of mismatched or corrupted modalities. Across voice, content, alignment, and clarity discrepancies, it preserves both accuracy and correlation better than baseline methods, confirming the strength of its correlation-aware architecture and contrastive learning in real-world multimodal scenarios with imperfectly aligned or noisy streams.
IV-H Ablation Studies
@@
For modality-specific feature extraction, we applied tailored approaches to address the distinct characteristics of three different data types. To gauge the effectiveness of these approaches, we specifically examined the transformation of the Mel spectrum matrix for audio data and the application of ResNet50 for image feature mapping. Table X quantitatively evaluates the effectiveness of different feature extraction strategies in the MCE module across various modalities. The strategies include the use of Mel spectrum transformation for audio data, ResNet50 for image data, both methods combined, and a baseline with no specialized feature extraction methods. Metrics assessed include loss values for audio ($Loss_{A}$), text ($Loss_{T}$), and visual ($Loss_{V}$) data, along with the average loss ($Loss_{ave}$) and epoch time ($t_{epoch}$). The results demonstrate the efficacy of modality-specific feature extraction. Tailored approaches such as Mel spectrum for audio and ResNet50 for visual data significantly reduce the losses across respective modalities, with the combined approach yielding the lowest average loss ($Loss_{ave}$ = 0.168) and shortest training time (41.02 seconds per epoch). This verifies the importance of employing specific feature extraction methods for each modality.
TABLE IX: The results of CorMulT while using different modality-specific feature extraction methods.
@@
We further evaluated CorMulT under different MCE pre-training configurations: no pre-training (None), audio and text only (A+T), visual and text only (V+T), and all three modalities (V+A+T). As shown in Table X, even without MCE pre-training, CorMulT still achieves competitive performance due to its robust MulT-based backbone. Nevertheless, incorporating additional pre-trained modalities within MCE steadily improves the results because the model acquires richer cross-modal correlation knowledge. In particular, expanding from a single pre-trained modality to two or three yields increasingly better synergy with CorMulT’s crossmodal attention layers, ultimately boosting sentiment analysis accuracy. These findings confirm that while the underlying MulT structure provides a strong starting point, the model’s performance benefits significantly from our correlation-aware design and modality pre-training.
TABLE X: The results of CorMulT while using different modalities for pre-training.